Binkz:

Kammu VedBraat sells plumbing equipment. He gets an annual salary of $17,496.00 plus 9.5% commission on all of his sales. If his sales for January was $2,460.00, what was his monthly for January?

Karla:

His salary for 12 months(1 year)is 17496$ So his salary for 1 month(January)=17496/12=1458 As written he would get commission on all his sale for January his sale=2460 Thus his commission for January is 9.5%*2460 =233.7 So his total income/salary for January= salary of January + commission in january he received Thus it's 1458+233.7=1691.7$

fwval:

e